# A Particle Algorithm for Mean-Field Variational Inference

### v2

@@ 5.3.5 Step 5: Establishing the recursive inequality
Part of the proof follows from the idea of Lemma 24 in Durmus and Moulines (2019). Define $V_{i}(\cdot)=\bar{V}_{i}(\cdot,q_{*}^{-i})$. Note that $Y_{t}^{i,j}\sim q_{*}^{i}$ and the density function of $q_{*}^{i}$ is proportional to $e^{-V_{i}(\cdot)}$. Then,
$\mathrm{d}Y_{t}^{i,j}=-V_{i}^{\prime}(Y_{t}^{i,j})\mathrm{d}t+\sqrt{2}\mathrm{d}B_{t}^{i,j},$
By Ito’s lemma,
$\mathrm{d}V_{i}^{\prime}(Y_{t}^{i,j})$ $=[-V_{i}^{\prime\prime}(Y_{t}^{i,j})V_{i}^{\prime}(Y_{t}^{i,j})+V_{i}^{\prime\prime\prime}(Y_{t}^{i,j})]\mathrm{d}t+\sqrt{2}V_{i}^{\prime\prime}(Y_{t}^{i,j})\mathrm{d}B_{t}^{i,j}.$
As a result,
$\bar{V}_{i}^{\prime}(Y_{t}^{i,j},q_{*}^{-i})-\bar{V}_{i}^{\prime}(Y_{0}^{i,j},q_{*}^{-i})=V_{i}^{\prime}(Y_{t}^{i,j})-V_{i}^{\prime}(Y_{0}^{i,j})$
$=\underbrace{\int_{0}^{t}[-V_{i}^{\prime\prime}(Y_{s}^{i,j})V_{i}^{\prime}(Y_{s}^{i,j})+V_{i}^{\prime\prime\prime}(Y_{s}^{i,j})]\mathrm{d}s}_{\alpha_{t}^{i,j}}+\sqrt{2}\underbrace{\int_{0}^{t}V_{i}^{\prime\prime}(Y_{s}^{i,j})\mathrm{d}B_{s}^{i,j}}_{\beta_{t}^{i,j}}.$
Let $\{\mathcal{F}_{t}\}_{t\geq 0}$ be the filtration generated by the process $\{(Y_{t}^{i,j},B_{t}^{i,j})\}_{t\geq 0}$ and the array $\bm{X}_{n}$. Since $\alpha_{t}^{i,j}$ is $\mathcal{F}_{t}$-adapted and $\beta_{t}^{i,j}$ is a $\mathcal{F}_{t}$-martingale, we have $\mathbb{E}(\beta_{t}^{i,j}|\mathcal{F}_{t})=0$ and $\mathbb{E}(\alpha_{t}^{i,j}\beta_{t}^{i,j})=0$.
Note that $A^{i,j}=\int_{0}^{h}(\alpha_{t}^{i,j}+\sqrt{2}\beta_{t}^{i,j})\mathrm{d}t$ and
$\mathbb{E}[(U_{+}^{i,j}-Y_{+}^{i,j})A^{i,j}]$ $=\mathbb{E}\bigg{(}(U_{+}^{i,j}-Y_{+}^{i,j})\int_{0}^{h}(\alpha_{t}^{i,j}+\sqrt{2}\beta_{t}^{i,j})\mathrm{d}t\bigg{)}$
$=\int_{0}^{h}\mathbb{E}[(U_{+}^{i,j}-Y_{+}^{i,j})(\alpha_{t}^{i,j}+\sqrt{2}\beta_{t}^{i,j})]\mathrm{d}t.$
Since $U_{+}^{i,j}-Y_{+}^{i,j}$ is $\mathcal{F}_{0}$-measurable, $\mathbb{E}[(U_{+}^{i,j}-Y_{+}^{i,j})\beta_{t}^{i,j}]=0$. We obtain that
$-2\mathbb{E}[(U_{+}^{i,j}-Y_{+}^{i,j})A^{i,j}]$ $=-2\int_{0}^{h}\mathbb{E}[(U_{+}^{i,j}-Y_{+}^{i,j})\alpha_{t}^{i,j}]\mathrm{d}t$
$\leq\int_{0}^{h}\Big{(}\varepsilon\mathbb{E}|U_{+}^{i,j}-Y_{+}^{i,j}|^{2}+\varepsilon^{-1}\mathbb{E}|\alpha_{t}^{i,j}|^{2}\Big{)}\mathrm{d}t$
$=\varepsilon h\mathbb{E}|U_{+}^{i,j}-Y_{+}^{i,j}|^{2}+\varepsilon^{-1}\int_{0}^{h}\mathbb{E}|\alpha_{t}^{i,j}|^{2}\mathrm{d}t,\qquad\forall\varepsilon>0.$
On the other hand, $A^{i,j}=\int_{0}^{h}(\alpha_{t}^{i,j}+\sqrt{2}\beta_{t}^{i,j})\mathrm{d}t$ and implies that
$\mathbb{E}|A^{i,j}|^{2}$ $=h^{2}\mathbb{E}\bigg{(}\frac{1}{h}\int_{0}^{h}(\alpha_{t}^{i,j}+\sqrt{2}\beta_{t}^{i,j})\mathrm{d}t\bigg{)}^{2}\leq h\int_{0}^{h}\mathbb{E}(\alpha_{t}^{i,j}+\sqrt{2}\beta_{t}^{i,j})^{2}\mathrm{d}t$
$=h\int_{0}^{h}(\mathbb{E}|\alpha_{t}^{i,j}|^{2}+2\mathbb{E}|\beta_{t}^{i,j}|^{2})\mathrm{d}t.$
@@
$-2\mathbb{E}[(U_{+}^{i,j}-Y_{+}^{i,j})A^{i,j}]+\mathbb{E}|A^{i,j}|^{2}$
$\leq\varepsilon h\mathbb{E}|U_{+}^{i,j}-Y_{+}^{i,j}|^{2}+(\varepsilon^{-1}+h)\int_{0}^{h}\mathbb{E}|\alpha_{t}^{i,j}|^{2}\mathrm{d}t+2h\int_{0}^{h}\mathbb{E}|\beta_{t}^{i,j}|^{2}\mathrm{d}t.$ (5.8)
We now work on $\mathbb{E}|\alpha_{t}^{i,j}|^{2}$ and $\mathbb{E}|\beta_{t}^{i,j}|^{2}$. By Ito isometry and Lemma 5.2, we have
$\mathbb{E}|\beta_{t}^{i,j}|^{2}=\mathbb{E}\bigg{|}\int_{0}^{t}V_{i}^{\prime\prime}(Y_{s}^{i,j})\mathrm{d}B_{s}^{i,j}\bigg{|}^{2}=\mathbb{E}\bigg{(}\int_{0}^{t}|V_{i}^{\prime\prime}(Y_{s}^{i,j})|^{2}\mathrm{d}s\bigg{)}\leq L^{2}t$
and $2h\int_{0}^{h}\mathbb{E}|\beta_{t}^{i,j}|^{2}\mathrm{d}t\leq L^{2}h^{3}$. In addition,
$|\alpha_{t}^{i,j}|^{2}$ $\leq\bigg{(}\int_{0}^{t}[L|V_{i}^{\prime}(Y_{s}^{i,j})|+\widetilde{L}]\mathrm{d}s\bigg{)}^{2}=\bigg{(}L\int_{0}^{t}|V_{i}^{\prime}(Y_{s}^{i,j})|\mathrm{d}t+\widetilde{L}t\bigg{)}^{2}$
$\leq 2L^{2}\bigg{(}\int_{0}^{t}|V_{i}^{\prime}(Y_{s}^{i,j})|\mathrm{d}t\bigg{)}^{2}+2\widetilde{L}^{2}t^{2}=2L^{2}t^{2}\bigg{(}\frac{1}{t}\int_{0}^{t}|V_{i}^{\prime}(Y_{s}^{i,j})|\mathrm{d}t\bigg{)}^{2}+2\widetilde{L}^{2}t^{2}$
$\leq 2L^{2}t^{2}\cdot\frac{1}{t}\int_{0}^{t}|V_{i}^{\prime}(Y_{s}^{i,j})|^{2}\mathrm{d}t+2\widetilde{L}^{2}t^{2}.$
Since $Y_{s}^{:,j}\sim q_{*}$, we have
$\mathbb{E}|V_{i}^{\prime}(Y_{s}^{i,j})|^{2}$ $=\mathbb{E}|\bar{V}_{i}^{\prime}(Y_{s}^{i,j},q_{*}^{-i})|^{2}=\mathbb{E}\Big{|}\mathbb{E}_{\bm{y}\sim q_{*}^{-i}}[\partial_{i}V(Y_{s}^{i,j},\bm{y})]\Big{|}^{2}$
$=\mathbb{E}\Big{|}\mathbb{E}[\partial_{i}V(\bm{Y}_{s}^{:,j})|Y_{s}^{i,j}]\Big{|}^{2}\leq\mathbb{E}|\partial_{i}V(\bm{Y}_{s}^{:,j})|^{2}=\mathbb{E}_{\bm{x}\sim q_{*}}|\partial_{i}V(\bm{x})|^{2}.$
and
$\mathbb{E}|\alpha_{t}^{i,j}|^{2}\leq 2t^{2}[L^{2}\mathbb{E}_{\bm{x}\sim q_{*}}|\partial_{i}V(\bm{x})|^{2}+\widetilde{L}^{2}].$
Hence,
$\int_{0}^{h}\mathbb{E}|\alpha_{t}^{i,j}|^{2}\mathrm{d}t\leq\frac{2h^{3}}{3}[L^{2}\mathbb{E}_{\bm{x}\sim q_{*}}|\partial_{i}V(\bm{x})|^{2}+\widetilde{L}^{2}].$
Finally, we obtain from (5.8) that
$-2\mathbb{E}[(U_{+}^{i,j}-Y_{+}^{i,j})A^{i,j}]+\mathbb{E}|A^{i,j}|^{2}$
$\leq\varepsilon h\mathbb{E}|U_{+}^{i,j}-Y_{+}^{i,j}|^{2}+(\varepsilon^{-1}+h)h^{3}[L^{2}\mathbb{E}_{\bm{x}\sim q_{*}}|\partial_{i}V(\bm{x})|^{2}+\widetilde{L}^{2}]+L^{2}h^{3}.$
5.5 Proof of Claim 5.2
Denote by $\mathcal{F}^{i}$ the $\sigma$-field generated by $(\bm{X}_{n},\bm{Y}^{i,:})$. Fix and $(i,j)\in[m]\times[N]$. $\bm{U}^{i,:}$ are $\mathcal{F}^{i}$-measurable. Fact 5.1 implies that conditioned on $\mathcal{F}^{i}$, we have $\bm{Y}^{-i,j}\sim q_{*}^{-i}$ and $\bm{U}^{-i,j}\sim q_{\bm{X}_{n}}^{-i}$. As a result,
$\mathbb{E}[\partial_{i}V(\bm{Y}^{:,j})|\mathcal{F}^{i}]=\mathbb{E}[\partial_{i}V(Y^{i,j},\bm{Y}^{-i,j})|\mathcal{F}^{i}]=\mathbb{E}_{\bm{y}\sim q_{*}^{-i}}\partial_{i}V(Y^{i,j},\bm{y})=\bar{V}_{i}^{\prime}(Y^{i,j},q_{*}^{-i}),$
$Y_{+}^{i,j}=Y^{i,j}-h\bar{V}_{i}^{\prime}(Y^{i,j},q_{*}^{-i})=\mathbb{E}[Y^{i,j}-h\partial_{i}V(\bm{Y}^{:,j})|\mathcal{F}^{i}].$
Similarly, we have $\mathbb{E}[\partial_{i}V(\bm{U}^{:,j})|\mathcal{F}^{i}]=\bar{V}_{i}^{\prime}(U^{i,j},q_{\bm{X}_{n}}^{-i})$ and
$U_{+}^{i,j}=U^{i,j}-h\bar{V}_{i}^{\prime}(U^{i,j},q_{\bm{X}_{n}}^{-i})=\mathbb{E}[U^{i,j}-h\partial_{i}V(\bm{U}^{:,j})|\mathcal{F}^{i}].$
Consequently,
$\mathbb{E}|U_{+}^{i,j}-Y_{+}^{i,j}|^{2}$ $=\mathbb{E}\Big{|}\mathbb{E}\Big{(}[U^{i,j}-h\partial_{i}V(\bm{U}^{:,j})]-[Y^{i,j}-h\partial_{i}V(\bm{Y}^{:,j})]\Big{|}\mathcal{F}^{i}\Big{)}\Big{|}^{2}$
$\leq\mathbb{E}\Big{|}[U^{i,j}-h\partial_{i}V(\bm{U}^{:,j})]-[Y^{i,j}-h\partial_{i}V(\bm{Y}^{:,j})]\Big{|}^{2}.$
@@
$\sum_{i=1}^{m}\mathbb{E}|U_{+}^{i,j}-Y_{+}^{i,j}|^{2}\leq\mathbb{E}\|\bm{\phi}(\bm{U}^{:,j})-\bm{\phi}(\bm{Y}^{:,j})\|_{2}^{2}\leq(1-\alpha h)^{2}\mathbb{E}\|\bm{U}^{:,j}-\bm{Y}^{:,j}\|_{2}^{2},$
where we used Lemma 6.2 and the assumption $0<h<2/(\alpha+L)$. Summing over $(i,j)$, we get
$\frac{1}{N}\sum_{(i,j)\in[m]\times[N]}\mathbb{E}|U_{+}^{i,j}-Y_{+}^{i,j}|^{2}\leq(1-\alpha h)^{2}\mathbb{E}\|\bm{U}-\bm{Y}\|_{\mathrm{F}}^{2}/N.$
@@ 6 Technical lemmas
Let $m\in\mathbb{Z}_{+}$ and $p_{i},q_{i}\in\mathcal{P}(\mathbb{R})$ for all $i\in[m]$. Then,
$\mathcal{W}_{2}^{2}\bigg{(}p_{1}\otimes\cdots\otimes p_{m},~{}q_{1}\otimes\cdots\otimes q_{m}\bigg{)}=\sum_{i=1}^{m}\mathcal{W}_{2}^{2}(p_{i},q_{i}).$
Proof of Lemma 6.1.
For any $i\in[m]$, let $\gamma_{i}$ be a $\mathcal{W}_{2}$-optimal coupling between $p_{i}$ and $q_{i}$. Then, $\gamma_{1}\otimes\cdots\otimes\gamma_{m}$ is a coupling between $p$ and $q$. For $(\bm{X},\bm{Y})\sim\gamma$, we have
$\mathcal{W}_{2}^{2}\bigg{(}p_{1}\otimes\cdots\otimes p_{m},~{}q_{1}\otimes\cdots\otimes q_{m}\bigg{)}\leq\mathbb{E}\|\bm{X}-\bm{Y}\|_{2}^{2}=\sum_{i=1}^{m}\mathbb{E}|X_{i}-Y_{i}|^{2}=\sum_{i=1}^{m}\mathcal{W}_{2}^{2}(p_{i},q_{i}).$
To prove the opposite direction of the bound, define $\gamma$ as a $\mathcal{W}_{2}$-optimal coupling between $p$ and $q$. If we draw $(\bm{X},\bm{Y})\sim\gamma$, then $X_{i}\sim p_{i}$, $Y_{i}\sim q_{i}$, and
$\mathcal{W}_{2}^{2}\bigg{(}p_{1}\otimes\cdots\otimes p_{m},~{}q_{1}\otimes\cdots\otimes q_{m}\bigg{)}=\mathbb{E}\|\bm{X}-\bm{Y}\|_{2}^{2}=\sum_{i=1}^{m}\mathbb{E}|X_{i}-Y_{i}|^{2}\geq\sum_{i=1}^{m}\mathcal{W}_{2}^{2}(p_{i},q_{i}).$
Lemma 6.2 (Lemma 1 in Dalalyan (2017a)).
Let Assumption 4.1 hold and $0<h<2/(\alpha+L)$. Define $\bm{\phi}(\bm{x})=\bm{x}-h\nabla V(\bm{x})$, $\forall\bm{x}\in\mathbb{R}^{mr}$. We have
$\|\bm{\phi}(\bm{x})-\bm{\phi}(\bm{y})\|_{2}\leq(1-\alpha h)\cdot\|\bm{x}-\bm{y}\|_{2},\qquad\forall\bm{x},\bm{y}\in\mathbb{R}^{mr}.$
@@
Under Assumption 4.1, we have $\mathbb{E}_{\bm{x}\sim q_{*}}[\nabla V(\bm{x})]=\bm{0}$, $\mathbb{E}_{\bm{x}\sim q_{*}}\|\nabla V(\bm{x})\|_{2}^{2}\leq mL^{2}/\alpha$, and $\operatorname{\rm cov}_{\bm{x}\sim q_{*}}(\bm{x})\preceq\alpha^{-1}\bm{I}_{m}$.
Proof of Lemma 6.3.
@@
Define $V_{1}(\cdot)=\bar{V}_{1}(\cdot,q_{*}^{-1})$. The density function of $q_{*}^{1}$ is proportional to $e^{-V_{1}(\cdot)}$, $V_{1}^{\prime\prime}(\cdot)\geq\alpha$ (Lemma 5.2), and thus $\lim_{x\to\pm\infty}V_{1}(x)=+\infty$. Hence,
@@
Based on the above, $\mathbb{E}_{\bm{x}\sim q_{*}}\nabla V(\bm{x})=\bm{0}$.
Consequently, $\mathbb{E}_{\bm{x}\sim q_{*}}|\partial_{1}V(\bm{x})|^{2}=\operatorname{{\rm var}}_{\bm{x}\sim q_{*}}[\partial_{1}V(\bm{x})]$. Note that $q_{*}$ is a product distribution with strongly log-concave marginals (Lemma 5.2). According to Bakry and Émery (2006), $q_{*}$ satisfies the Poincaré inequality
$\operatorname{{\rm var}}_{\bm{x}\sim q_{*}}f(\bm{x})\leq\alpha^{-1}\mathbb{E}_{\bm{x}\sim q_{*}}\|\nabla f(\bm{x})\|_{2}^{2},\qquad\forall f\in C^{1}(\mathbb{R}^{m}).$ (6.1)
@@
Finally, choose any unit vector $\bm{u}\in\mathbb{R}^{m}$. By (6.1), we have
@@
Therefore, $\operatorname{\rm cov}_{x\sim q_{*}^{i}}(\bm{x})\leq\alpha^{-1}\bm{I}_{m}$. ∎
Lemma 6.4 (Corollary 6.12 in Bobkov and Ledoux (2019)).
@@
Let Assumption 4.1 hold and $N\geq 2$. Consider the notations in Definition 3.2, and define $\bm{Y}\in\mathbb{R}^{m\times N}$ with $\{\bm{Y}^{:,j}\}_{j=1}^{N}$ being i.i.d. samples from $q_{*}$. There exists a universal constant $c$ such that
$\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{Y}},q_{*})\leq\frac{cm\log N}{\alpha N}.$
Proof of Corollary 6.1.
By Lemma 6.1,
$\mathcal{W}_{2}^{2}(q_{\bm{Y}},q_{*})=\sum_{i=1}^{m}\mathcal{W}_{2}^{2}(q_{\bm{Y}}^{i},q_{*}^{i})=\mathcal{W}_{2}^{2}\bigg{(}\frac{1}{N}\sum_{j=1}^{N}\delta_{Y^{i,j}},q_{*}^{i}\bigg{)}.$
By Lemmas 6.3 and 6.4, there exists a universal constant $c$ such that
$\mathbb{E}\mathcal{W}_{2}^{2}\bigg{(}\frac{1}{N}\sum_{j=1}^{N}\delta_{Y^{i,j}},q_{*}^{i}\bigg{)}\leq\frac{c\log N}{\alpha N},\qquad\forall i\in[m],$
and thus,
$\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{Y}},q_{*})\leq\frac{cm\log N}{\alpha N}.$

### v3

@@ 5.5.5 Step 5: Proving the recursive inequality
Let $\{\sigma_{i}\}_{i=1}^{m}$ be i.i.d. permutations of $[N]$ that are independent of $(\bm{X}_{n},\bm{Y}_{nh})$. Denote by $\mathcal{F}^{i}$ the $\sigma$-field generated by $(\bm{X}_{n}^{i,:},\bm{Y}_{nh}^{i,:},\sigma_{i})$. Define two arrays $\bm{U},\bm{Z}\in\mathbb{R}^{m\times N}$ through $U^{i,j}=X_{n}^{i,\sigma_{i}(j)}$ and $Z^{i,j}=Y_{nh}^{i,\sigma_{i}\circ\tau_{n,i}(j)}$. We have
$\sum_{j=1}^{N}|U_{+}^{i,j}-Y_{+}^{i,j}|^{2}$ $=\sum_{j=1}^{N}\{[X_{n}^{i,j}-h\bar{V}_{i}^{\prime}(X_{n}^{i,j},q_{\bm{X}_{n}}^{-i})]-[Y_{nh}^{i,\tau_{n,i}(j)}-h\bar{V}_{i}^{\prime}(Y_{nh}^{i,\tau_{n,i}(j)},q_{*}^{-i})]\}^{2}$
$=\sum_{j=1}^{N}\{[U^{i,j}-h\bar{V}_{i}^{\prime}(U^{i,j},q_{\bm{X}_{n}}^{-i})]-[Z^{i,j}-h\bar{V}_{i}^{\prime}(Z^{i,j},q_{*}^{-i})]\}^{2}$
$=\sum_{j=1}^{N}\{[U^{i,j}-h\bar{V}_{i}^{\prime}(U^{i,j},q_{\bm{U}_{n}}^{-i})]-[Z^{i,j}-h\bar{V}_{i}^{\prime}(Z^{i,j},q_{*}^{-i})]\}^{2},$ (5.17)
where the last equality follows from the fact that $q_{\bm{X}_{n}}=q_{\bm{U}}$.
By definition, $\bm{U}^{i,:}$ and $\bm{Z}^{i,:}$ are $\mathcal{F}^{i}$-measurable. Conditioned on $\mathcal{F}^{i}$, we have $\bm{Z}_{nh}^{-i,j}\sim q_{*}^{-i}$ and $\bm{U}^{-i,j}\sim q_{\bm{U}}^{-i}$. As a result,
$\mathbb{E}[\partial_{i}V(\bm{Z}^{:,j})|\mathcal{F}^{i}]=\mathbb{E}[\partial_{i}V(Z^{i,j},\bm{Z}^{-i,j})|\mathcal{F}^{i}]=\mathbb{E}_{\bm{z}\sim q_{*}^{-i}}\partial_{i}V(Z^{i,j},\bm{z})=\bar{V}_{i}^{\prime}(Z^{i,j},q_{*}^{-i}),$
$Z^{i,j}-h\bar{V}_{i}^{\prime}(Z^{i,j},q_{*}^{-i})=\mathbb{E}[Z^{i,j}-h\partial_{i}V(\bm{Z}^{:,j})|\mathcal{F}^{i}].$
Similarly, we have $\mathbb{E}[\partial_{i}V(\bm{U}^{:,j})|\mathcal{F}^{i}]=\bar{V}_{i}^{\prime}(U^{i,j},q_{\bm{U}}^{-i})$ and
$U^{i,j}-h\bar{V}_{i}^{\prime}(U^{i,j},q_{\bm{U}}^{-i})=\mathbb{E}[U^{i,j}-h\partial_{i}V(\bm{U}^{:,j})|\mathcal{F}^{i}].$
@@
$\mathbb{E}\{[U^{i,j}-h\bar{V}_{i}^{\prime}(U^{i,j},q_{\bm{U}_{n}}^{-i})]-[Z^{i,j}-h\bar{V}_{i}^{\prime}(Z^{i,j},q_{*}^{-i})]\}^{2}$
$=\mathbb{E}\Big|\mathbb{E}\Big([U^{i,j}-h\partial_{i}V(\bm{U}^{:,j})]-[Z^{i,j}-h\partial_{i}V(\bm{Z}^{:,j})]\Big|\mathcal{F}^{i}\Big)\Big|^{2}$
$\leq\mathbb{E}\Big|[U^{i,j}-h\partial_{i}V(\bm{U}^{:,j})]-[Z^{i,j}-h\partial_{i}V(\bm{Z}^{:,j})]\Big|^{2}.$
@@
$\sum_{i=1}^{m}\mathbb{E}\{[U^{i,j}-h\bar{V}_{i}^{\prime}(U^{i,j},q_{\bm{U}_{n}}^{-i})]-[Z^{i,j}-h\bar{V}_{i}^{\prime}(Z^{i,j},q_{*}^{-i})]\}^{2}$
$\leq\mathbb{E}\|\bm{\phi}(\bm{U}^{:,j})-\bm{\phi}(\bm{Z}^{:,j})\|_{2}^{2}\leq(1-\alpha h)^{2}\mathbb{E}\|\bm{U}^{:,j}-\bm{Z}^{:,j}\|_{2}^{2},$
where we used Lemma˜6.3 and the assumption $0<h<2/(\alpha+L)$. Based on the above, we use (5.17) to get
$\frac{1}{N}\sum_{(i,j)\in[m]\times[N]}\mathbb{E}|U_{+}^{i,j}-Y_{+}^{i,j}|^{2}$ $\leq\frac{(1-\alpha h)^{2}}{N}\sum_{(i,j)\in[m]\times[N]}\mathbb{E}|U^{i,j}-Z^{i,j}|^{2}$
$=\frac{(1-\alpha h)^{2}}{N}\sum_{(i,j)\in[m]\times[N]}\mathbb{E}|X_{n}^{i,j}-Y_{nh}^{i,\tau_{n,i}(j)}|^{2}.$
@@ 6 Technical lemmas
Lemma 6.1.
Let $\mu$, $\nu$ and $\rho$ be random probability distributions on $\mathbb{R}^{m}$. Then,
$\Big|\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(\mu,\nu)}-\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(\mu,\rho)}\Big|\leq\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(\nu,\rho)}.$
Proof of Lemma˜6.1.
By the triangle inequality of $\mathcal{W}_{2}$, we have $|\mathcal{W}_{2}(\mu,\nu)-\mathcal{W}_{2}(\mu,\rho)|\leq\mathcal{W}_{2}(\nu,\rho)$. By Minkowski’s inequality,
$\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(\nu,\rho)}$ $\geq\sqrt{\mathbb{E}|\mathcal{W}_{2}(\mu,\nu)-\mathcal{W}_{2}(\mu,\rho)|^{2}}\geq\Big|\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(\mu,\nu)}-\sqrt{\mathbb{E}\mathcal{W}_{2}^{2}(\mu,\rho)}\Big|\,.$
@@
Let $\{d_{i}\}_{i=1}^{m}$ be positive integers, and $p_{i},q_{i}\in\mathcal{P}(\mathbb{R}^{d_{i}})$ for $i\in[m]$. We have
$\mathcal{W}_{2}^{2}(p_{1}\otimes\cdots\otimes p_{m},~q_{1}\otimes\cdots\otimes q_{m})=\sum_{i=1}^{m}\mathcal{W}_{2}^{2}(p_{i},q_{i}).$
Proof of Lemma˜6.2.
Let $p=p_{1}\otimes\cdots\otimes p_{m}$ and $q=q_{1}\otimes\cdots\otimes q_{m}$. For $i\in[m]$ and $\bm{x}\in\mathbb{R}^{d_{1}+\cdots+d_{m}}$, denote by $\bm{x}_{i}$ the $(1+\sum_{j=1}^{i-1}d_{j})$-th through $\sum_{j=1}^{i}d_{j}$-th coordinates of $\bm{x}$; let $\gamma_{i}$ be a $\mathcal{W}_{2}$-optimal coupling between $p_{i}$ and $q_{i}$. Then, $\gamma_{1}\otimes\cdots\otimes\gamma_{m}$ is a coupling between $p$ and $q$. For $(\bm{X},\bm{Y})\sim\gamma$, we have
$\mathcal{W}_{2}^{2}(p,q)\leq\mathbb{E}\|\bm{X}-\bm{Y}\|_{2}^{2}=\sum_{i=1}^{m}\mathbb{E}\|\bm{X}_{i}-\bm{Y}_{i}\|^{2}=\sum_{i=1}^{m}\mathcal{W}_{2}^{2}(p_{i},q_{i}).$
To prove the opposite direction, define $\gamma$ as a $\mathcal{W}_{2}$-optimal coupling between $p$ and $q$. If we draw $(\bm{X},\bm{Y})\sim\gamma$, then $\bm{X}_{i}\sim p_{i}$, $\bm{Y}_{i}\sim q_{i}$, and
$\mathcal{W}_{2}^{2}(p,q)=\mathbb{E}\|\bm{X}-\bm{Y}\|_{2}^{2}=\sum_{i=1}^{m}\mathbb{E}\|\bm{X}_{i}-\bm{Y}_{i}\|^{2}\geq\sum_{i=1}^{m}\mathcal{W}_{2}^{2}(p_{i},q_{i}).$
Lemma 6.3 (Lemma 1 in Dal17).
Let Assumption 4.1 hold and $0<h<2/(\alpha+L)$. Define $\bm{\phi}(\bm{x})=\bm{x}-h\nabla V(\bm{x})$, $\forall\bm{x}\in\mathbb{R}^{m}$. We have
$\|\bm{\phi}(\bm{x})-\bm{\phi}(\bm{y})\|_{2}\leq(1-\alpha h)\cdot\|\bm{x}-\bm{y}\|_{2},\qquad\forall\bm{x},\bm{y}\in\mathbb{R}^{m}.$
@@
Let Assumption 4.1 hold. We have
$\mathbb{E}_{\bm{x}\sim q_{*}}[\nabla V(\bm{x})]=\boldsymbol{0},\qquad\mathbb{E}_{\bm{x}\sim q_{*}}\|\nabla V(\bm{x})\|_{2}^{2}\leq mL^{2}/\alpha,\qquad\operatorname{\rm cov}_{\bm{x}\sim q_{*}}(\bm{x})\preceq\alpha^{-1}\bm{I}_{m}.$
For any $\mu\in\mathcal{P}(\mathbb{R}^{m-1})$ and $x\in\mathbb{R}$, we have $\alpha\leq\bar{V}_{i}^{\prime\prime}(x,\mu)\leq L$.
Proof of Lemma˜6.4.
@@
Define $V_{1}(\cdot)=\bar{V}_{1}(\cdot,q_{*}^{-1})$. The density function of $q_{*}^{1}$ is proportional to $e^{-V_{1}(\cdot)}$, $V_{1}^{\prime\prime}(\cdot)\geq\alpha$ (Lemma˜6.4), and thus $\lim_{x\to\pm\infty}V_{1}(x)=+\infty$. Hence,
@@
Based on the above, $\mathbb{E}_{\bm{x}\sim q_{*}}[\nabla V(\bm{x})]=\boldsymbol{0}$.
Consequently, $\mathbb{E}_{\bm{x}\sim q_{*}}|\partial_{1}V(\bm{x})|^{2}=\operatorname{{\rm var}}_{\bm{x}\sim q_{*}}[\partial_{1}V(\bm{x})]$. Note that $q_{*}$ is a product distribution with strongly log-concave marginals (Lemma˜6.4). According to BEm06, $q_{*}$ satisfies the Poincaré inequality
$\operatorname{{\rm var}}_{\bm{x}\sim q_{*}}[f(\bm{x})]\leq\alpha^{-1}\mathbb{E}_{\bm{x}\sim q_{*}}\|\nabla f(\bm{x})\|_{2}^{2},\qquad\forall f\in C^{1}(\mathbb{R}^{m}).$ (6.1)
@@
Next, choose any unit vector $\bm{u}\in\mathbb{R}^{m}$. By (6.1), we have
@@
Therefore, $\operatorname{\rm cov}_{\bm{x}\sim q_{*}^{i}}(\bm{x})\leq\alpha^{-1}\bm{I}_{m}$. Finally, the proof of $\alpha\leq\bar{V}_{i}^{\prime\prime}(x,\mu)\leq L$ is immediate and thus omitted. ∎
Lemma 6.5 (Corollary 6.12 in BLe19).
@@
Let Assumption 4.1 hold, $N\geq 2$, and $\bm{X}\in\mathbb{R}^{m\times N}$ be a random array. Suppose that the columns of $\bm{X}$ are drawn independently from a product distribution $q$, whose marginals are log-concave with variance bounded by $\sigma^{2}$. There exists a universal constant $c$ such that
$\mathbb{E}\mathcal{W}_{2}^{2}(q_{\bm{X}},q)\leq\frac{cm\sigma^{2}\log N}{N},$
where $q_{\bm{X}}$ is constructed according to Definition˜3.2.